15 houses line Indian Creek. The average space between each house exceeds the average width of each house by 80 feet. The sidewalk, 2920 feet in length, begins at a point 30 feet before the first house and ends at a point 30 feet beyond the last house. What is the average width, in feet, of each house?
Answer by ankor@dixie-net.com(22740)

15 houses line Indian Creek.
The average space between each house exceeds the average width of each house by 80 feet.
The sidewalk, 2920 feet in length, begins at a point 30 feet before the first house and ends at a point 30 feet beyond the last house.
What is the average width, in feet, of each house?
:
Let w = the av width of the houses
:
Subtract the 30 ft from either side of the block of houses
2920 - 60 = 2860
:
Total width of the houses = 15w
Total spaces between houses = 14(w+80)
:
15w + 14(w+80) = 2860
15w + 14w + 1120 = 2860
29w = 2860 - 1120
29w = 1740
w = 1740/29
w = 60 ft the av width of each house